Circular No. 7640

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001cj IN UGC 8399 M. Ganeshalingam and W. D. Li, University of California at Berkeley,

report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on May 30.3 (mag ∼ 15.9) and 31.3 UT (mag ∼ 15.7). The new object is located at α = 13^{h}21^{m}44^{s}.85, δ = +31^{o}14′48′′.6 (equinox 2000.0), which is 7′′.6 west and 35′′.2 north of the nucleus of UGC 8399. A KAIT image of the same field taken on May 17.3 showed nothing at this position (limiting mag ∼ 18.5).

L. Wang, Lawrence Berkeley National Laboratory, writes: “SN 2001cj was observed using the Kitt Peak 2.1-m telescope (+ GCAM) on June 1. A fully reduced spectrum shows that it is of type Ia and is before optical maximum. The blueshift of the asbsorption trough of the P-Cyg profile of the Si II 635.5-nm line is at 10 580 km/s after being corrected for the redshift of the host galaxy.”

NO SUPERNOVA 2001bn M. Phillips, Carnegie Observatories; and K. Krisciunas, Cerro Tololo

Interamerican Observatory, report that a spectrum obtained with the Du Pont 2.5-m telescope at Las Campanas Observatory on May 27.17 UT shows that SN 2001bn (cf. IAUC 7624) is not a supernova, but rather a blue variable star with strong, narrow H I Balmer absorption lines at z ∼ 0.0.

COMET 29P/SCHWASSMANN-WACHMANN 1 This comet is undergoing another outburst, as shown by the following

total magnitudes (CCD unless otherwise indicated): Apr. 22.80 UT, 15.5 (A. Nakamura, Kuma, Ehime, Japan, 0.60-m reflector); 28.44, 15.7 (A. Hale, Cloudcroft, NM, 0.2-m reflector); May 17.69, 13.2 (K. Kadota, Ageo, Saitama, Japan, 0.18-m reflector); 18.71, 13.4 (Kadota); 19.73, 12.0 (K. Yoshimoto, Hirao, Yamaguchi, Japan, 0.25-m reflector; visual); 27.77, 13.5 (Nakamura).

SUPERNOVA 2001V IN NGC 3987 CCD R magnitudes by K. Hornoch, Lelekovice, Czech Republic (0.35-

m reflector): Apr. 24.951 UT, 15.8; 27.908, 15.9; 29.879, 15.9; May 9.845, 15.9; 10.863, 16.0; 11.956, 16.0; 13.927, 16.1; 20.900, 16.2; 24.917, 16.3.

2001 June 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green